# ThinkingViT: Matryoshka Thinking Vision Transformer for Elastic Inference

### v2

@@ Abstract
Vision Transformers (ViTs) deliver SOTA performance, yet their fixed computational budget prevents scalable deployment across heterogeneous hardware. Recent Matryoshka-style Transformer architectures mitigate this by embedding nested subnetworks within a single model to enable scalable inference. However, these models allocate the same amount of compute to all inputs, regardless of their complexity, which leads to inefficiencies. To address this, we introduce ThinkingViT, a nested ViT architecture that employs progressive thinking stages to dynamically adjust inference computation based on input difficulty. ThinkingViT first activates a small subset of the most important attention heads to produce an initial prediction. If the prediction confidence exceeds a predefined threshold, inference terminates early. Otherwise, within the same backbone, it activates a larger subset of attention heads and conducts a new forward pass. This process continues iteratively until the model reaches the predefined confidence level or exhausts its maximum capacity. To boost the performance of subsequent rounds, we introduce a Token Recycling approach that fuses the input embeddings with the embeddings from the previous stage. Experiments show that ThinkingViT surpasses nested baselines by up to 2.0 percentage points (p.p.) in accuracy at the same throughput and by up to 2.9 p.p. at equal GMACs on ImageNet-1K. We show that the backbone-preserving design of ThinkingViT allows it to serve as a plug-in upgrade for ViTs in downstream tasks such as semantic segmentation. We also demonstrate that ThinkingViT transfers effectively to other architectures such as Swin.
@@
Figure 1: Comparison of ThinkingViT with MatFormer(NeurIPS’24) [devvrit2024matformer], HydraViT(NeurIPS’24) [haberer2024hydravit], SortedNet(NeurIPS’23-W) [valipour2023sortednet], and DynaBERT(NeurIPS’20) [hou2020dynabert], evaluated in terms of GMACs and throughput on an A100. All baselines have a dynamic width within a standard ViT backbone and are trained following the training recipes in [touvron2021training]. ThinkingViT is trained with two progressive thinking stages using 3 and 6 heads, and consistently surpasses baseline by up to 2.0 p.p. at the same throughput and by up to 2.9 p.p. at the same GMACs. See Appendix H for a comparison with smaller baseline models.
Figure 2: Nested progressive inference with Token Recycling in ThinkingViT. After embedding the input, ThinkingViT first activates a subset of the model, including the first attention heads (e.g., 50%), to produce an initial prediction. Due to the training procedure, these heads capture the most important features. If the certainty exceeds a threshold (easy inputs), inference terminates early to save computation. Otherwise, the resulting tokens are fused back into the input via a learnable scaling factor $\alpha$, which controls how much prior knowledge is recycled. The model then thinks more by reprocessing the fused tokens using a larger subset of the attention heads (e.g., 100%) for a refined prediction. ThinkingViT enables elastic inference across different hardware budgets by adjusting the confidence threshold. The number of thinking stages and the proportion of attention heads activated at each stage can be flexibly configured based on efficiency and accuracy trade-offs, see Section 4.2.
@@ 1 Introduction
Motivation: Vision Transformer (ViT) [dosovitskiy2021an] have achieved state-of-the-art results across numerous image recognition tasks [liu2022swin, hatamizadeh2024mambavision], yet their non-elastic inference pipelines impose a uniform and often excessive computational cost. This lack of flexibility poses a significant limitation in real-world deployments, where devices vary widely in computational power and latency constraints [xu2025edgevit]. From high-throughput servers to mobile and embedded platforms, modern applications require elastic inference [cai2024flextron], where a model can adjust its computational footprint to match the capabilities of the target hardware. Without such adaptability, ViTs struggle to meet the performance and efficiency tradeoffs necessary for scalable and widespread deployment.
@@
ThinkingViT: Inspired by recent advances in thinking-based Large Language Modelss (LLMs) [guo2025deepseek, shao2024deepseekmath], we introduce ThinkingViT, a ViT built upon the vanilla ViT architecture [dosovitskiy2021an] that dynamically adjusts its inference effort based on input complexity through nested progressive thinking stages. ThinkingViT begins by activating a subset of the nested network, including the most important attention heads (e.g., the top 50%) to generate an initial prediction along with a certainty score. If the certainty is sufficiently high (the "aha" moment [guo2025deepseek]), inference terminates early. Otherwise, the processed embeddings are fused with the original input embeddings and passed through a larger subset of the nested network, using an increased set of attention heads (e.g., 100%) to perform a more thorough re-evaluation. This token fusion mechanism allows the model to avoid thinking from scratch; instead, it recycles the knowledge gained in the first pass, refining its prediction with improved accuracy; see Figure 2. Our proposed progressive thinking mechanism enables the model not only to think more, but also to think more powerfully when processing ambiguous inputs, ensuring that more challenging examples receive greater representational capacity, i.e., more attention heads. In this pipeline, the prediction certainty serves as the central mechanism that governs elastic inference: the model halts early for cases with high certainty and progressively deepens computation for more uncertain inputs. By adjusting the certainty threshold, the model naturally balances performance against efficiency, supporting elastic inference without requiring any separate routing mechanism. Additionally, nesting all stages within a single unified model avoids parameter duplication and enables more efficient training and improved accuracy.
@@ 2 Related Work
Nested Models: Slimmable networks first demonstrated that a single model can operate at multiple widths using shared weights and width-specific normalization [yu2018slimmable, yu2019universally]. Many subsequent works build on this idea. For instance, MatFormer [devvrit2024matformer], based on Matryoshka Representation Learning [kusupati2022matryoshka], introduces multiple nested subsets within the hidden layer of MLP. DynaBERT [hou2020dynabert] slices the Multi-head Attention (MHA) and MLP layers, but does not slice embeddings across layers since it relies on knowledge distillation. SortedNet [valipour2023sortednet] generalizes nesting across MLP s, Normalization Layer (NORM), MHA, and embedding dimensions, although it retains a fixed number of attention heads. HydraViT [haberer2024hydravit] and Slicing ViT [zhang2024slicing] enable slicing across embeddings, NORM, MLP, and MHA, and support a dynamic number of heads. However, like the other mentioned methods, they apply a fixed compute budget per input, limiting their ability to adapt based on input complexity and leading to inefficiencies under constrained resources.
@@
Thinking: Recent advancements in LLM have introduced mechanisms for adaptive reasoning depth [el2025competitive, jaech2024openai]. These methods often use reinforcement learning to dynamically adjust the number of reasoning steps based on input complexity [guo2025deepseek, shao2024deepseekmath]. However, excessive reasoning length can inflate inference cost without proportional gains in accuracy [kumar2025overthink]. Motivated by this, recent efforts seek to preserve the benefits of deep reasoning while avoiding redundant computation by introducing test-time adaptability or confidence-based early stopping [muennighoff2025s1, zeng2025revisiting]. These models, however, reuse the same network in each round, which, as reported in Table 1, results in limited gains. In contrast, ThinkingViT scales up its capacity during rethinking to overcome this challenge.
@@ 3.1 Nested Vision Transformer - Sorting Heads
ThinkingViT is built on top of the standard ViT architecture. Let $V_{D,H}$ be a ViT model with $H$ attention heads and the embedding dimension of $D$. ViT begins by tokenizing the input image $x$ into $P$ non-overlapping patches, each of which is projected into a $D$-dimensional embedding vector $\mathcal{E}^{D}$ using a CNN-based patch embedding function. After adding positional encodings, the resulting sequence is processed by $L$ standard Transformer blocks:
$\quad z^{l}=\text{Block}^{l}(z^{l-1})\quad\text{for }l=1,\dots,L$ (1)
@@ 3.2 Looping nested Vision Transformers through Token Recycling
After constructing the nested architecture, ThinkingViT begins by using the smallest subnetwork $V_{d_{i},h_{i}}$, with the embedding dimension of $d_{i}$ and $h_{i}$ attention heads. This subnetwork processes the input image to predict the class and generate token embeddings $z^{L}$. Subsequently, to refine its prediction, the model expands computational capacity by activating a larger subnetwork $V_{d_{j},h_{j}}(x)$, featuring $h_{j}$ attention heads, where $h_{i}<h_{j}$ and $d_{i}<d_{j}$. Afterwards, it fuses the produced token embeddings $z^{L}$ with the new input embeddings $\mathcal{E}^{d_{j}}(x)\in\mathbb{R}^{P\times d_{j}}$, through a projection layer and scaled by a learnable parameter $\alpha$, determining the weight of embeddings "recalled" from the previous step:
$\mathcal{E}^{d_{j}}_{\mathrm{fused}}=\alpha\cdot\mathrm{Proj}_{d_{i}\rightarrow d_{j}}(z^{L})+\mathcal{E}^{d_{j}}(x),$ (4)
This Token Recycling mechanism allows the model to reuse prior representations instead of reprocessing from scratch. The process continues iteratively and, unlike language models that reuse the same network for reasoning [guo2025deepseek], our approach increases model capacity by activating progressively larger subsets of attention heads, enhancing representational capacity at each step. Progressive expansion is essential in vision tasks, where naive recursion quickly plateaus, as shown in Table 1. Appendix E and Appendix D compare recycling and fusion strategies in detail.
@@ 3.5 Elastic inference
ThinkingViT supports elastic inference by using its own certainty to determine when to stop. The entropy threshold $\tau$ governs the trade-off between efficiency and accuracy. A low threshold enforces stricter confidence requirements, causing most inputs to undergo more thinking steps, which increases computational cost but improves accuracy. In contrast, a high threshold allows earlier exits for more inputs, reducing latency and computation cost at the expense of potential accuracy loss. This flexibility enables users to tune the model at runtime according to specific resource and performance needs.
Figure 3: Accuracy vs. GMACs for ThinkingViT variants on ImageNet-1K. 3H $\rightarrow$ 6H achieves the best trade-off, while 2H $\rightarrow$ 3H $\rightarrow$ 6H covers the widest range with only a slight accuracy drop. See Appendix K for details.
@@ 4 Evaluation
In this section, we analyze how entropy guides ThinkingViT’s adaptive computation (Sec. 4.2), evaluate it on ImageNet-1K and its variants against SOTA nested baselines (Sec. 4.3), extend the evaluation to hierarchical such as Swin Transformer (Sec. 4.6) and segmentation architectures (Sec. 4.5), compare with early-exit baselines (Sec. 4.7), and finally assess robustness on uniformly challenging datasets (Sec. 4.8).
@@ 4.1 Setup
Implementation details: We train ThinkingViT on ImageNet-1K [ILSVRC15] at a resolution of $224\times 224$, using models implemented in timm [rw2019timm] and following the setup of touvron2021training, with all models initialized with a pretrained DeiT-Tiny checkpoint. We train on a cluster with NVIDIA H100 GPUs, taking about 10 minutes per epoch on 2 GPUs. During prototyping, we conducted roughly 100 training runs, each lasting 300 epochs.
Baselines: We compare ThinkingViT against several width expansion baselines: MatFormer [devvrit2024matformer], which slices only the hidden layer of MLP while keeping MHA and embedding dimensions fixed; DynaBERT [hou2020dynabert], which slices both MHA and the hidden layer of MLP while keeping embeddings fixed; SortedNet [valipour2023sortednet], which slices all embeddings, NORM, MHA, and MLP, while keeping the number of heads fixed; and HydraViT [haberer2024hydravit], which slices MHA, embeddings, NORM, and MLP, while also changing the number of heads. Similar to HydraViT, ThinkingViT adopts a slicing strategy across the number of heads, MHA, embeddings, NORM, and MLP layers. However, unlike the above methods that follow static inference paths regardless of input difficulty, ThinkingViT introduces input-adaptive computation. In Section 4.7, we present comparisons to depth-expansion baselines.
Throughput and GMACs: We measure throughput and computational cost on a single NVIDIA A100 GPU with a batch size of 512. To evaluate throughput, we run the model under different confidence thresholds $\tau$. For each threshold, we record the per-image inference time across the entire validation set and compute the average throughput as the total number of images processed divided by the total inference time. By varying the threshold $\tau$, we obtain the throughput curves reported in the plots. For further details, see Appendix J. For GMACs, we follow a similar procedure. The total compute per image for threshold $\tau$ is then averaged over all samples as
@@ 4.2 Trade-offs in attention expansion
ThinkingViT expands computational capacity through progressive activation of attention heads. Generally, the number of thinking stages and the attention head expansion step size are hyperparameters, and their optimal configuration depends on the dataset and target efficiency objectives. In Figure 3, we evaluate this trade-off and find that the 3H $\rightarrow$ 6H configuration offers the most favorable balance between accuracy and compute on ImageNet-1K. The 2H $\rightarrow$ 3H $\rightarrow$ 6H variant spans the widest GMAC range among the high-performing models, while incurring only a small drop in accuracy compared to 3H $\rightarrow$ 6H. At the upper end, 3H $\rightarrow$ 6H $\rightarrow$ 12H achieves the highest final accuracy, surpassing 3H $\rightarrow$ 6H by 0.91 percentage points (p.p.), but raises the compute cost from 5.85 GMACs to 23.41 GMACs for only a marginal gain. This substantial increase in cost leads to lower overall efficiency compared to the 3H $\rightarrow$ 6H setup, underscoring the importance of aligning ThinkingViT’s thinking strategy with specific deployment goals. Since we focus on configurations that offer the best trade-off between accuracy and GMACs, we adopt 3H $\rightarrow$ 6H for subsequent experiments. For additional experiments and a larger depiction of Figure 3, see Appendix K.
@@
Figure 7: Load distribution of ThinkingViT 3H $\rightarrow$ 6H across entropy thresholds. Bars indicate usage share; numbers show accuracy (green: correct, red: incorrect) on ImageNet-1K. High 3H accuracy at low entropy confirms the effectiveness of entropy-based routing.
@@
Figure 8: (a): ThinkingViT with a Swin backbone outperforms standard Swin models, matching Swin-Base accuracy with far fewer GMACs. (b): Comparison of ThinkingViT with Early-Exit baselines (with Swin backbones), showing that iterative refinement provides improvement beyond early termination strategies.

### v3

@@ Abstract
Vision Transformers (ViTs) deliver SOTA performance, yet their fixed computational budget prevents scalable deployment across heterogeneous hardware. Recent Matryoshka-style Transformer architectures mitigate this by embedding nested subnetworks within a single model to enable scalable inference. However, these models allocate the same amount of compute to all inputs, regardless of their complexity, which leads to inefficiencies. To address this, we introduce ThinkingViT, a nested ViT architecture that employs progressive thinking stages to dynamically adjust inference computation based on input difficulty. ThinkingViT first activates a small subset of the most important attention heads to produce an initial prediction. If the prediction confidence exceeds a predefined threshold, inference terminates early. Otherwise, within the same backbone, it activates a larger subset of attention heads and conducts a new forward pass. This process continues iteratively until the model reaches the predefined confidence level or exhausts its maximum capacity. To boost the performance of subsequent rounds, we introduce a Token Recycling approach that fuses the input embeddings with the embeddings from the previous stage. Experiments show that ThinkingViT surpasses nested baselines by up to 2.0 percentage points (p.p.) in accuracy at the same throughput and by up to 2.9 p.p. at equal GMACs on ImageNet-1K. We show that the backbone-preserving design of ThinkingViT allows it to serve as a plug-in upgrade for ViTs in downstream tasks such as semantic segmentation. We also demonstrate that ThinkingViT transfers effectively to other architectures such as Swin Transformers.
@@
Figure 1: Comparison of ThinkingViT with MatFormer(NeurIPS’24) [devvrit2024matformer], HydraViT(NeurIPS’24) [haberer2024hydravit], SortedNet(NeurIPS’23-W) [valipour2023sortednet], and DynaBERT(NeurIPS’20) [hou2020dynabert], evaluated in terms of GMACs and throughput on an A100. All baselines have a dynamic width within a standard ViT backbone and are trained following the training recipes in [touvron2021training]. ThinkingViT is trained with two progressive thinking stages using 3 and 6 heads, and consistently surpasses baselines by up to 2.0 p.p. at the same throughput and by up to 2.9 p.p. at the same GMACs. See Appendix H for a comparison with smaller baseline models.
Figure 2: Nested progressive inference with Token Recycling in ThinkingViT. After embedding the input, ThinkingViT first activates a subset of the model, including the first attention heads (e.g., 50%), to produce an initial prediction. Due to the training procedure, these heads capture the most important features. If the certainty exceeds a threshold (easy inputs), inference terminates early to save computation. Otherwise, the resulting tokens are fused back into the input via a projection and a learnable scaling factor $\alpha$, which controls how much prior knowledge is recycled. The model then thinks more by reprocessing the fused tokens using a larger subset of the attention heads (e.g., 100%) for a refined prediction. ThinkingViT enables elastic inference across different hardware budgets by adjusting the confidence threshold. The number of thinking stages and the proportion of attention heads activated at each stage can be flexibly configured based on efficiency and accuracy trade-offs, see Section 4.2.
@@ 1 Introduction
Motivation: Vision Transformers (ViTs) [dosovitskiy2021an] have achieved state-of-the-art results across numerous image recognition tasks [liu2022swin, hatamizadeh2024mambavision], yet their non-elastic inference pipelines impose a uniform and often excessive computational cost. This lack of flexibility poses a significant limitation in real-world deployments, where devices vary widely in computational power and latency constraints [xu2025edgevit]. From high-throughput servers to mobile and embedded platforms, modern applications require elastic inference [cai2024flextron], where a model can adjust its computational footprint to match the capabilities of the target hardware. Without such adaptability, ViTs struggle to meet the performance and efficiency tradeoffs necessary for scalable and widespread deployment.
@@
ThinkingViT: Inspired by recent advances in thinking-based Large Language Models (LLMs) [guo2025deepseek, shao2024deepseekmath], we introduce ThinkingViT, a ViT built upon the vanilla ViT architecture [dosovitskiy2021an] that dynamically adjusts its inference effort based on input complexity through nested progressive thinking stages. ThinkingViT begins by activating a subset of the nested network, including the most important attention heads (e.g., the top 50%) to generate an initial prediction along with a certainty score. If the certainty is sufficiently high (the "aha" moment [guo2025deepseek]), inference terminates early. Otherwise, the processed embeddings are fused with the original input embeddings and passed through a larger subset of the nested network, using an increased set of attention heads (e.g., 100%) to perform a more thorough re-evaluation. This token fusion mechanism allows the model to avoid thinking from scratch; instead, it recycles the knowledge gained in the first pass, refining its prediction with improved accuracy; see Figure 2. Our proposed progressive thinking mechanism enables the model not only to think more, but also to think more powerfully when processing ambiguous inputs, ensuring that more challenging examples receive greater representational capacity, i.e., more attention heads. In this pipeline, the prediction certainty serves as the central mechanism that governs elastic inference: the model halts early for cases with high certainty and progressively deepens computation for more uncertain inputs. By adjusting the certainty threshold, the model naturally balances performance against efficiency, supporting elastic inference without requiring any separate routing mechanism. Additionally, nesting all stages within a single unified model avoids parameter duplication and enables more efficient training and improved accuracy.
@@ 2 Related Work
Nested Models: Slimmable networks first demonstrated that a single model can operate at multiple widths using shared weights and width-specific normalization [yu2018slimmable, yu2019universally]. Many subsequent works build on this idea. For instance, MatFormer [devvrit2024matformer], based on Matryoshka Representation Learning [kusupati2022matryoshka], introduces multiple nested subsets within the hidden layer of MLP. DynaBERT [hou2020dynabert] slices the Multi-head Attention (MHA) and MLP layers, but does not slice embeddings across layers since it relies on knowledge distillation. SortedNet [valipour2023sortednet] generalizes nesting across MLP, Normalization Layer (NORM), MHA, and embeddings, although it retains a fixed number of attention heads. HydraViT [haberer2024hydravit] and Slicing ViT [zhang2024slicing] enable slicing across embeddings, NORM, MLP, and MHA, and support a dynamic number of heads. However, like the other mentioned methods, they apply a fixed compute budget per input, limiting their ability to adapt based on input complexity and leading to inefficiencies under constrained resources.
@@
Thinking: LLMs have introduced mechanisms for adaptive reasoning depth [el2025competitive, jaech2024openai]. These methods often use reinforcement learning to dynamically adjust the number of reasoning steps based on input complexity [guo2025deepseek, shao2024deepseekmath]. However, excessive reasoning length can inflate inference cost without proportional gains in accuracy [kumar2025overthink]. Motivated by this, recent efforts seek to preserve the benefits of deep reasoning while avoiding redundant computation by introducing test-time adaptability or confidence-based early stopping [muennighoff2025s1, zeng2025revisiting]. These models, however, reuse the same network in each round, which, as reported in Table 1, results in limited gains. In contrast, ThinkingViT scales up its capacity during rethinking to overcome this challenge.
@@ 3.1 Nested Vision Transformer - Sorting Heads
ThinkingViT is built on top of the standard ViT architecture. Let $V_{D,H}$ be a ViT model with $H$ attention heads and the embedding dimension of $D$. ViT begins by tokenizing the input image $x$ into $P$ non-overlapping patches, each of which is projected into a $D$-dimensional embedding vector $\mathcal{E}^{D}$ using a convolution. After adding positional encodings, the resulting sequence is processed by $L$ Transformer blocks, where $z_{0}=\mathcal{E}^{D}$:
$\quad z_{l}=\text{Block}_{l}(z_{l-1})\quad\text{for }l=1,\dots,L$ (1)
@@ 3.2 Looping nested Vision Transformers through Token Recycling
After constructing the nested architecture, ThinkingViT begins by using a small subnetwork $V_{d_{i},h_{i}}$, with the embedding dimension of $d_{i}$ and $h_{i}$ attention heads. This subnetwork processes the input image to predict the class and generate token embeddings $z_{L}$. Subsequently, to refine its prediction, the model expands computational capacity by activating a larger subnetwork $V_{d_{j},h_{j}}(x)$, featuring $h_{j}$ attention heads, where $h_{i}<h_{j}$ and $d_{i}<d_{j}$. Afterwards, it fuses the produced token embeddings $z_{L}$ with the new input embeddings $\mathcal{E}^{d_{j}}(x)\in\mathbb{R}^{P\times d_{j}}$, through a projection layer and scaled by a learnable parameter $\alpha$, determining the weight of embeddings "recalled" from the previous step:
$\mathcal{E}^{d_{j}}_{\mathrm{fused}}=\alpha\cdot\mathrm{Proj}_{d_{i}\rightarrow d_{j}}(z_{L})+\mathcal{E}^{d_{j}}(x),$ (4)
This Token Recycling mechanism allows the model to reuse prior representations instead of reprocessing from scratch. The process continues iteratively and, unlike language models that reuse the same network for reasoning [guo2025deepseek], our approach increases model capacity by activating progressively larger subsets of attention heads, enhancing representational capacity at each step. Progressive expansion is essential in vision tasks, where naïve recursion quickly plateaus, as shown in Table 1. Appendix E and Appendix D compare recycling and fusion strategies in detail.
@@ 3.5 Elastic inference
ThinkingViT supports elastic inference by using its own certainty to determine when to stop. The entropy threshold $\tau$ governs the trade-off between efficiency and accuracy. A low threshold enforces stricter confidence requirements, causing most inputs to undergo more thinking steps, which increases computational cost but improves accuracy. In contrast, a high threshold allows more inputs to exit early, reducing latency and computation cost at the expense of potential accuracy loss. This flexibility enables users to tune the model at runtime according to specific resource and performance needs.
Figure 3: Accuracy vs. GMACs for ThinkingViT variants on ImageNet-1K. 3H $\rightarrow$ 6H achieves the best trade-off, while 2H $\rightarrow$ 3H $\rightarrow$ 6H covers the widest range with only a slight accuracy drop. See Appendix L for details.
@@ 4 Evaluation
In this section, we analyze how entropy guides ThinkingViT’s adaptive computation (Sec. 4.2), evaluate it on ImageNet-1K and its variants against SOTA nested baselines (Sec. 4.3), extend the evaluation to segmentation architectures (Sec. 4.5) and hierarchical architectures such as Swin Transformer (Sec. 4.6), compare with early-exit baselines (Sec. 4.7), and finally assess robustness on uniformly challenging datasets (Sec. 4.8).
@@ 4.1 Setup
Implementation details: We train ThinkingViT on ImageNet-1K [ILSVRC15] at a resolution of $224\times 224$, using models implemented in timm [rw2019timm] and following the setup of touvron2021training, with all models initialized with a pretrained DeiT-Tiny checkpoint. We train on a cluster with NVIDIA H100 GPUs, taking about 10 minutes per epoch on 2 GPUs.
Baselines: We compare ThinkingViT against several width expansion baselines: MatFormer [devvrit2024matformer], which slices only the hidden layer of the MLP while keeping MHA and embedding dimensions fixed; DynaBERT [hou2020dynabert], which slices both MHA and the hidden layer of MLP while keeping embeddings fixed; SortedNet [valipour2023sortednet], which slices all embeddings, NORM, MHA, and MLP, while keeping the number of heads fixed; and HydraViT [haberer2024hydravit], which slices MHA, embeddings, NORM, and MLP, while also changing the number of heads. Similar to HydraViT, ThinkingViT adopts a slicing strategy across the number of heads, MHA, embeddings, NORM, and MLP layers. However, unlike the above methods that follow static inference paths regardless of input difficulty, ThinkingViT introduces input-adaptive computation. In Section 4.7, we present comparisons to depth-expansion baselines.
Throughput and GMACs: We measure throughput and computational cost on a single NVIDIA A100 GPU with a batch size of 512. To evaluate throughput, we run the model under different confidence thresholds $\tau$. For each threshold, we record the per-image inference time across the entire validation set and compute the average throughput as the total number of images processed divided by the total inference time. By varying the threshold $\tau$, we obtain the throughput curves reported in the plots. For further details, see Appendix K. For GMACs, we follow a similar procedure. The total compute per image for threshold $\tau$ is then averaged over all samples as
@@ 4.2 Trade-offs in attention expansion
ThinkingViT expands computational capacity through progressive activation of attention heads. Generally, the number of thinking stages and the attention head expansion step size are hyperparameters, and their optimal configuration depends on the dataset and target efficiency objectives. In Figure 3, we evaluate this trade-off and find that the 3H $\rightarrow$ 6H configuration offers the most favorable balance between accuracy and compute on ImageNet-1K. The 2H $\rightarrow$ 3H $\rightarrow$ 6H variant spans the widest GMAC range among the high-performing models, while incurring only a small drop in accuracy compared to 3H $\rightarrow$ 6H. At the upper end, 3H $\rightarrow$ 6H $\rightarrow$ 12H achieves the highest final accuracy, surpassing 3H $\rightarrow$ 6H by 0.91 percentage points (p.p.), but raises the compute cost from 5.85 GMACs to 23.41 GMACs for only a marginal gain. This substantial increase in cost leads to lower overall efficiency compared to the 3H $\rightarrow$ 6H setup, underscoring the importance of aligning ThinkingViT’s thinking strategy with specific deployment goals. Since we focus on configurations that offer the best trade-off between accuracy and GMACs, we adopt 3H $\rightarrow$ 6H for subsequent experiments. For additional experiments and a larger depiction of Figure 3, see Appendix L.
@@
Figure 7: Load distribution of ThinkingViT 3H $\rightarrow$ 6H across entropy thresholds. Bars indicate load distribution; numbers show accuracy (green: correct, red: incorrect) on ImageNet-1K. High 3H accuracy at low entropy confirms the effectiveness of entropy-based routing.
@@
Figure 8: (a): ThinkingViT with a Swin backbone outperforms standard Swin models, matching Swin-Base accuracy with far fewer GMACs. (b): Comparison of ThinkingViT with Early-Exit baselines (with Swin backbones), showing that iterative expansion provides improvement beyond early termination strategies.